# JEL: C81, O32

### v2

@@ 3.3.1. Prompt Design
We identify three general prompt formats, which differ both in the amount of detail provided about our classification task and in the structure of the requested model completion. The simplest prompt provides a version of our sample definition, 3.1, and the text of an abstract, and asks the model to return ‘TRUE’ if the abstract satisfies these criteria. Otherwise, it returns ‘FALSE’. The text of this prompt is displayed in Figure E.1. We refer to this prompt as Prompt 1.0. A second, more complicated prompt provides the same definition, with a set of examples of publication characteristics that do and do not satisfy the definition. Here, the prompt asks the model to return either ‘TRUE’ or the name of a specific excluded category. The text of this prompt is displayed in Figure E.7. We refer to this prompt as Prompt 2.0. A third, more complex prompt provides the same definitions and examples, but asks the model to return either ‘TRUE’ or an explanation of why the record does or does not satisfy our sample definition. The text of this prompt is displayed in Figure E.10. We refer to this prompt as Prompt 3.0. We devise initial language for each prompt iteratively, using small numbers of records from our hand-labelled validation data.
@@
Table 3.1 reports estimates of the true positive rate and false positive rate for each prompt, in both models, computed in the validation sample. Observe—for example, with Prompts 1.1, 1.2, and 1.3 queried to GPT-4—that even small changes in the text of a prompt yield substantial differences in performance. We attempt a second round of prompt iteration (Prompt 1.3). We observe that performance deteriorates with even small modification. Thus, we select the two highest performing prompts—one for each proprietary model. We use Prompt 2.0 to extract weak labels using GPT-3.5, and Prompt 1.2 to extract weak labels using GPT-4. Figure 3.3 displays estimates of the true and false positive rates computed in the test data, for both proprietary models, with black dots. These prompts are used to extract 64,000 labels—TRUE/FALSE flags—using GPT-3.5 and GPT-4. We refer to these as our set of “noisy labels.”
@@
Figure 3.3. Receiver Operating Characteristic
Notes: Figure 3.3 displays estimates of the receiver operating characteristic of several models used to classify the publications indexed by PubMed / MEDLINE according to whether they satisfy the restrictions enumerated in 3.1. All metrics are computed in the testing split of the hand-labelled publications. Estimates of the true positive rate and false positive rate of OpenAI’s proprietary models GPT-3 and GPT-4 are indicated with black dots. The curves give the performance of four fine-tuned, open-source large language models, in addition to an ensemble model. The blue dots indicate the true positive rate and false positive rate of the models used to construct the “Conservative,” “Moderate,” and “Liberal” samples of clinical trials.
GPT-4 performs remarkably well. For every 71 true positives, GPT-4 identifies 29 false positives, while achieving a true positive rate of approximately 0.974. This represents a substantial improvement over existing methods. However, practical and substantive considerations mitigate the applicability of proprietary models for classification of the complete baseline sample. Practically, it is—and likely will remain—prohibitively expensive to deploy GPT-4 at this scale. [^35] We incur a cost of roughly \$4,500 to extract noisy labels for 64,000 abstracts. Extrapolating this figure to the full sample of 1.8 million abstracts gives a price of approximately \$130,000. Substantively, proprietary models are black boxes. Their details and substance are not public and are known to change at regular intervals.
@@ 3.3.2. Fine-Tuning
We compute noisy labels for 64,000 randomly selected publications in our baseline sample using the best performing prompts for both GPT-3 and GPT-4. These noisy labels are used as data to train off-the-shelf language models from two architecture classes: (1) RoBERTa (Liu et al.,, 2019) and (2) BioMedBERT (Gu et al.,, 2021). Both models are available in either “baseline” or “large” varieties. [^36] There are many open-source large language models (e.g., LLaMa, Mistral, Pythia). Many are several orders of magnitude (e.g., 7-70 billion parameters) larger than RoBERTA ($\sim$ 350 million parameters). We selected RoBERTA and BioMedBERT via trial-and-error that included testing the performance of these larger models. In our setting, RoBERTA outperforms LLaMA (both 7 and 70 billion parameter versions, trained with QLoRA). Both models exhibited comparable performance, but RoBERTa was substantially faster and simpler to train. The BioMedBERT model is an instance of a BERT model (Devlin et al.,, 2019), that has itself been fine-tuned on the abstracts of the publications indexed in PubMed. We systematically vary model architecture, size, and pre-training regimen. We observe no large differences in performance stemming from these hyper-parameters. Sections C.4 and C.5 give further details concerning the fine-tuning process and the results of these experiments, respectively.
@@ 3.4. Performance
Given the text of an abstract, our fine-tuned language models output a probability that the publication satisfies the restrictions enumerated in 3.1. Publications whose probabilities fall above a chosen threshold are classified as belonging in our sample. Figure 3.3 displays estimates of the true and false positive rates, computed with the test data, as we vary this threshold for both sizes of the RoBERTa and BioMedBERT models. [^37] Each of the models whose results are displayed in Figure 3.3 are trained with noisy labels obtained with GPT-4. We document in Section C.5 that models trained with noisy labels obtained from GPT-3.5 or with the 1000 hand-labeled records in the training data perform significantly worse. The figure additionally displays the performance of an ensemble model estimated in the training data. This model is obtained by estimating a logistic regression of the hand-labels on the probabilities output by all four models displayed Figure 3.3. The ensemble model is used to produce our final sample.
@@
The fine-tuned models are able to match the performance of GPT-4 in the test data. We choose three thresholds according to the stringency with which they enforce the sample restrictions. The estimated true and false positive rates associated with these points are displayed in Figure 3.3 and labeled as “Conservative,” “Moderate,” and “Liberal.” Our preferred sample is associated with the conservative threshold. [^38] The liberal model may be particularly useful for conducting literature reviews, where a near-perfect true positive rate is needed. For every 82 true positives, the conservative model identifies 18 false positives. [^39] In the test data, the conservative model assigns incorrect labels to 27 of 993 papers. We conduct an error analysis. See Section C.6 for further details. In 13 cases, there is a clear error. In 14 cases, however, errors are associated with records that were difficult to categorize for a human labeller. For example, PubMed record 32737793 is flagged as satisfying 3.1 with all three thresholds, but is a literature review. By contrast, PubMed record 27880726 was categorized as satisfying 3.1 twice by a human labeller. It is excluded from both the conservative and moderate model-generated samples. On inspection, the abstract does not explicitly state that the study enrolled human subjects, but hints that it may have been conducted in an animal model. Review of the associated full text confirms that this study did, in fact, enroll only rats. We report results associated with the moderate and liberal thresholds in the appendix as tests of robustness. Our final, conservative, sample consists of 152,027 publications classified as satisfying the restrictions enumerated in 3.1.
@@
Each table displayed in Appendix A compares the contents of our sample to the counts of potential clinical trials plotted in Figure 3.1. Although certain elements of PubMed metadata—including the union of all NLM tags—capture many of the records in our final sample, we confirm that they miss many records that we flag as trials and include many records that do not satisfy 3.1. No combination of existing search strategies, then, classifies records with the same accuracy or precision as our final sample.
4. Disentangling Trends in Medical Research
@@
Figure 3.1 indicates that, across measures, the total quantity of medical research appears to be increasing, including since 2010.404040We find similar patterns using alternative cuts of records indexed in PubMed. The total number of unique authors listed on scientific papers, the total number of citations given and received by scientific papers, the total number of scientific journals, and the total number of pages in scientific journals all exhibit similar patterns. These patterns are consistent with documented increases in the quantity of scientific papers in Bloom et al., (2020) and Park et al., (2023), as well as economy-wide increases in research effort reported in Goldin et al., (2024).
In this section, we use our census of clinical trials to decompose this trend. We contrast trends in the production of clinical trials with those documented for other forms of medical research. Differences between these series are informative in two ways. First, existing estimates of the quantity of medical research, such as those plotted in Figure 3.1, capture some combination of these two series (e.g., Bloom et al.,, 2020; Park et al.,, 2023; Chu and Evans,, 2021). Second, any proxy for the quantity of medical research is difficult to interpret if its relationship to the underlying research changes over time. This decomposition highlights the limitations of empirical strategies that treat counts of scientific publications as a proxy for research quantity.
@@
There are many potential ways to construct samples of other (i.e., non-clinical trial) medical research. Our preferred approach defines this category as the collection of publications that cite clinical trials. (We confirm in Section 4.3.1 that this set includes only medical research.) Conceptually, of course, this “citing” sample is a subset of non-clinical trial medical research. Alternative methods of constructing a measure of “other” research introduce additional, complicating conceptual problems. Approaches that rely on keyword searches, NLM tags, etc., as Section 3 and its appendices make clear, capture different sets of records over time, as a consequence of changes in database structure. Approaches that examine other publications in the same journals as clinical trials will mechanically suggest shifts in total quantities, if the number of pages in journals change over time or if certain journals become online-only. [^41] Ioannidis et al., (2023) document that the number of medical journals is increasing. In 2024, the Annals of Emergency Medicine, the highest impact journal in the medical sub-field of emergency medicine, shifted to an online-only format. Brainard, (2020) discusses dozens of instances in which medical journals transitioned to online-only formats, then ceased publication entirely.
@@
4.1. Decomposing Types of Medical Research
Figure 4.1 documents that the quantity of clinical trials—the black line—has remained essentially constant over the past ten years. In contrast, the quantity of other medical research—the colored lines—has increased substantially. Observe that we display the percent growth in the quantity of clinical trials in our sample published in each calendar year, alongside the growth in the set of papers that cite clinical trials, constructed in various ways to account for the truncation discussed above. Each series is normalized to begin at zero. This presentation masks substantial differences in the levels of these series. There are 10,903 clinical trials in our census in 2010, and 30,841 publications that cite clinical trials (using three-year citations) in 2013.
@@ 4.2. Clinical Trials as a Productivity Indicator
Recall the very simple conceptual framework from Section 2. We assume, following e.g., Bloom et al., (2020), that the left-hand side of the “idea production function” (2.1) is constant in this setting. If clinical trials, as captured in our data, are a stable proxy for inputs $S_{t}$—in the sense that published trials capture the same amount of underlying research in each period—then an unchanging number of clinical trials implies that research productivity is constant. We return to the appropriate interpretation of rising trends in other medical research below. Here, we collect evidence that clinical trials are a stable proxy for inputs.
@@
These proxy variables capture slightly different conceptual objects, which are helpful in understanding different aspects of the production function (2.1). Stability in public funding suggests minimal changes, at least in the cross section, in allocation of resources to clinical trials over time. [^44] NIH RePORTER data allow us to tab total expenditures associated with grants acknowledged in these publications. These grant-based expenditure measures, too, are stable. Note, however, that linking research grant dollars to specific scientific papers is an imprecise exercise. See Li, (2017) for an extended discussion of this measurement challenge. Stability in citation patterns, in contrast, suggests that the “usefulness” of information produced by trials has been stable across this period.
Figure 4.3 provides a more granular view of this heterogeneity. We plot quantiles of the distribution of the number of (three-year) citations to publications in our census from publications in leading journals, for each year between 2010 and 2019.454545Appendix D presents these results using alternative cuts of our data and five-year citations. The patterns remain unchanged. Grey areas at the bottom of each plot represent publications that receive zero citations. This distribution is stable over our time period. Observe that we elongate the y-axis to highlight the right-tail of the citation distribution. The quantiles corresponding to the right tail of the citation distribution are constant across years. The left-tail is similarly stable: the share of records receiving zero citations is essentially constant, year-on-year.
@@
We interpret this set of facts as evidence that our measured quantity of clinical trials captures an object that, though heterogeneous, is stable over time. Thus, suggestively, research productivity is similarly stable. We cannot reject that there have been shifts in upstream, basic inputs into clinical trials or changes in factor prices during this period, which might suggest that the underlying quantity of resources directed toward research is, in fact, changing over time. In our view, however, there is little rigorous evidence that supports such an argument. A small number of papers examine frictions in markets for basic research. See, for example, Hill and Stein, (2021) and Myers, (2020). To our knowledge, this literature has not provided evidence of substantial shifts in inputs or factor prices over time, which in turn would increase the cost or decrease the productivity of basic science. [^46] Anecdotal evidence suggests that there may have been substantial decreases in the cost of basic science during this time period, which would suggest the opposite conclusion in our simple framework—that research might be more productive over this period. See Scannell et al., (2012). On factor prices, as discussed in Section 3, nearly all estimates of research and development costs are self-reported by the pharmaceutical industry. Any measured increases in these expenditures that cannot be independently verified should be viewed with caution, especially in light of our findings of stability. [^47] Herper, (2022) collects anecdotal evidence, which has been widely cited, on sharply rising costs of clinical trials. Although the magnitudes included in this discussion are striking, they capture the impressions of a small number of industry stakeholders and, to our knowledge, have not been verified in a more detailed analysis.
For clinical trials to be an informative proxy for the quantity of medical research—and for the black line in Figure 4.1, then, to capture the relevant measure of inputs—we must make two additional assumptions. First, the existence of publication bias means that scientific publications do not perfectly capture research inputs (Andrews and Kasy,, 2019). In our setting—where the object of interest is the trend in research over time—we need only make a relatively weak assumption about the extent of publication bias: we assume that changes in publication bias, year-on-year, do not perfectly offset changes in the quantity of actual research.
@@
Second, we must assume that the number of publications associated with each clinical trial is constant across our period of interest (i.e., that the amount of research “embedded” in each publication is not systematically changing over time). We view this assumption as reasonable for three reasons. First, by 2010, nearly all scientific journals in our sample had adopted rules requiring authors of clinical trial publications to adhere to a standardized format when reporting trial results (Schulz et al.,, 2010). Under these rules, the structure of a paper and its contents are pre-determined; there is little scope for authors to, for example, split results across several publications. Second, when constructing our sample of clinical trials, we have aimed to exclude those publications that re-analyze data generated by an existing study. Third, more suggestively, we use data on clinical trial publications supported by National Institutes of Health grants to examine the quantity of grant dollars per reported clinical trial publication over this period. We find no evidence of a change in this relationship between funding and publication during our period of interest. [^48] We inspect this relationship by collecting records from the NIH RePORTER database, for years 2010 to 2019. Specifically, we collect records of the total award associated with funded projects and the quantity of any publications flagged by our procedure as clinical trials. For the median NIH grant, we find that an award of roughly \$100,000 per clinical trial, and a mean of roughly \$200,000 per trial. This is stable across years of data.

### v3

@@ 2.4.1. Prompt Design
We identify three general prompt formats, which differ both in the amount of detail provided about our classification task and in the structure of the requested model completion. The simplest prompt provides a version of our sample definition, 2.1, and the text of an abstract, and asks the model to return ‘TRUE’ if the abstract satisfies these criteria. Otherwise, it returns ‘FALSE’. The text of this prompt is displayed in Figure E.1. We refer to this prompt as Prompt 1.0. A second, more complicated prompt provides the same definition, with a set of examples of publication characteristics that do and do not satisfy the definition. Here, the prompt asks the model to return either ‘TRUE’ or the name of a specific excluded category. The text of this prompt is displayed in Figure E.7. We refer to this prompt as Prompt 2.0. A third, more complex prompt provides the same definitions and examples, but asks the model to return either ‘TRUE’ or an explanation of why the record does or does not satisfy our sample definition. The text of this prompt is displayed in Figure E.10. We refer to this prompt as Prompt 3.0. We devise initial language for each prompt iteratively, using a small number of records from our hand-labelled validation data.
@@
Table 2.1 reports estimates of the true positive rate and false positive rate for each prompt, in both models, computed in the validation sample. Observe—for example, with Prompts 1.1, 1.2, and 1.3 queried to GPT-4—that even small changes in the text of a prompt yield substantial differences in performance. We attempt a second round of prompt iteration (Prompt 1.3). We observe that performance deteriorates with even small modification. Thus, we select the two highest performing prompts—one for each proprietary model. We use Prompt 2.0 to extract weak labels using GPT-3.5, and Prompt 1.2 to extract weak labels using GPT-4.
@@
The performance of GPT-4 represents a substantial improvement over existing methods. However, practical and substantive considerations mitigate the applicability of proprietary models for classification of the complete baseline sample. Practically, it is—and likely will remain—prohibitively expensive, both computationally and financially, to deploy GPT-4 at this scale. [^21] We incur a cost of roughly \$4,500 to extract noisy labels for 64,000 abstracts. Extrapolating this figure to the full sample of 1.8 million abstracts gives a price of approximately \$130,000. Substantively, proprietary models are black boxes. Their details and substance are not public and are known to change at regular intervals.
@@ 2.4.2. Fine-Tuning
We compute noisy labels for a moderate number of randomly selected publications in our baseline sample using the best performing prompts for both GPT-3.5 and GPT-4. These noisy labels are used as data to train off-the-shelf BERT models from two architecture classes: (1) BigBird (Zaheer et al.,, 2021) and (2) BioMedBERT (Gu et al.,, 2021). [^22] There are many open-source large language models (e.g., LLaMa, Mistral, Pythia). Many are several orders of magnitude (e.g., 7-70 billion parameters) larger than BERT ($\sim$ 350 million parameters). We selected BigBird and BioMedBERT via trial-and-error that included testing the performance of these larger models, including LLaMA (both 7 and 70 billion parameter versions, trained with QLoRA). Both models exhibited comparable performance, but BigBird was substantially faster and simpler to train. Both models use medium-scale Transformer architectures, i.e., between 100 and 300 million parameters, pre-trained with the masked-language modeling (MLM) objective (Vaswani et al.,, 2023; Devlin et al.,, 2019). The models differ in their architectural details and pre-training corpora.
BigBird uses a sparse attention mechanism to reduce the computational cost of processing long text sequences. [^23] The computational complexity of regular attention is quadratic in sequence length, while BigBird reduces this to linear complexity. This enables efficient handling of long documents up to 4,096 tokens. In contrast, standard BERT models can only process 512 tokens. BigBird is useful because abstracts of clinical trials regularly exceed 512 tokens. Prior to pre-training, the model was warm-started from the RoBERTa checkpoint (Liu et al.,, 2019) and then pre-trained on the standard BERT corpus.242424 The standard BERT corpus consists of the Books (Zhu et al.,, 2015), CC-News (Guu et al.,, 2020), Stories (Trinh and Le,, 2019), and Wikipedia (Wikimedia Foundation,, 2023) datasets. We evaluate two model sizes: BigBird Base (125 million parameters) and BigBird Large (355 million parameters).
BioMedBert uses a domain-specific pre-training corpus sourced from articles on PubMed (Gu et al.,, 2021). It uses the same model architecture as RoBERTa, but a different, domain-specific token vocabulary optimized for PubMed articles. Like BigBird, we evaluate two model sizes: BioMedBert Base (125 million parameters) and BioMedBert Large (355 million parameters).
We fine-tune the pre-trained architectures to classify abstracts according to the inclusion and exclusion criteria enumerated in 2.1. We replace the language modeling classification head with a binary classification head consisting of a single linear layer and softmax activation. We fine-tune the full model (i.e. not just the classification head) with cross-entropy loss. [^25] We use the Adam optimizer with learning rate $1\times 10^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$ (determined via hyperparameter sweep). We use a maximum sequence length of 4096 tokens. If an abstract does not fit into the maximum sequence length provided by a model, the abstract is truncated. All models were trained on a single $V100$ GPU with 16GB of HBM.
@@ 2.5. Performance
Given the text of an abstract, our fine-tuned language models output a probability that the publication satisfies the restrictions enumerated in 2.1. Publications whose probabilities fall above a chosen threshold are classified as belonging in our sample. Figure 2.3 displays estimates of true and false positive rates, computed with the test data, as we vary this threshold for labels assigned by fine-tuned version of the base BigBird model. The top and bottom panels are trained on noisy labels extracted from GPT-3.5 and GPT-4, respectively. We vary the quantity of noisy labels used to train the language model. Additionally, in the bottom panel, we display the performance of a fine-tuned model trained with the 1000 hand-labelled observations in the training dataset.
Figure 2.3. Receiver Operating Characteristic by Data Size and Origin
Notes: Figure 2.3 displays estimates of the the receiver operating characteristic of several fine-tuned language models, based on the BigBird architecture. The models differ in terms of the source of their training data. Solid lines display the performance of models trained on noisy labels extracted from either GPT-3.5 or GPT-4. The dotted line displays the performance of a model trained on the hand-labeled data in the training dataset. Estimates of the true positive rate and false positive rate of GPT-3.5 and GPT-4 are indicated with black dots and are measured using the testing data.
Fine-tuned models trained with noisy labels extracted from GPT-3.5 outperform GPT-3.5. By contrast, fine-tuned models trained with noisy labels extracted from GPT-4 match the performance of GPT-4, and significantly outperform models trained with labels extracted from GPT-3.5 or with hand-labels. There appears to be a threshold where performance dramatically improves at around 8,000 training labels.
Figure 2.4. Receiver Operating Characteristic by Model Architecture
Notes: Figure 2.4 displays estimates of the receiver operating characteristic of several models used to classify the publications indexed by PubMed / MEDLINE according to whether they satisfy the restrictions enumerated in 2.1. All metrics are computed in the testing split of the hand-labelled publications. Estimates of the true positive rate and false positive rate of OpenAI’s proprietary models GPT-3 and GPT-4 are indicated with black dots. The curves give the performance of four fine-tuned, open-source large language models, in addition to an ensemble model. The blue dots indicate the true positive rate and false positive rate of the models used to construct the “Conservative,” “Moderate,” and “Liberal” samples of clinical trials.
Figure 2.4 displays analogous estimates for both sizes of the BigBird and BioMedBERT models. Both models are trained with 64,000 training labels extracted with GPT-4. Again, the fine-tuned models are able to match the performance of GPT-4 in the test data. The figure additionally displays the performance of an ensemble model estimated in the training data. This model is obtained by estimating a logistic regression of the hand-labels on the probabilities output by all four models displayed Figure 2.4. The ensemble model is used to produce our final sample.
@@
We choose three thresholds according to the stringency with which they enforce the sample restrictions. The estimated true and false positive rates associated with these points are displayed in Figure 2.4 and labeled as “Conservative,” “Moderate,” and “Liberal.” Our preferred sample is associated with the conservative threshold. [^26] The liberal model may be particularly useful for conducting literature reviews, where a near-perfect true positive rate is needed. For every 82 true positives, the conservative model identifies 18 false positives. [^27] In the test data, the conservative model assigns incorrect labels to 27 of 993 papers. We conduct an error analysis. See Section C.4 for further details. In 13 cases, there is a clear error. In 14 cases, however, errors are associated with records that were difficult to categorize for a human labeller. For example, PubMed record 32737793 is flagged as satisfying 2.1 with all three thresholds, but is a literature review. By contrast, PubMed record 27880726 was categorized as satisfying 2.1 twice by a human labeller. It is excluded from both the conservative and moderate model-generated samples. On inspection, the abstract does not explicitly state that the study enrolled human subjects, but hints that it may have been conducted in an animal model. Review of the associated full text confirms that this study did, in fact, enroll only rats. We report results associated with the moderate and liberal thresholds in the appendix as tests of robustness. Our final, conservative, sample consists of 152,027 publications classified as satisfying the sample restrictions.
@@
Each table displayed in Appendix A compares the contents of our sample to the counts of potential clinical trials plotted in Figure 2.1. Although certain elements of PubMed metadata—including the union of all NLM tags—capture many of the records in our final sample, we confirm that they miss many records that we flag as trials and include many records that do not satisfy 2.1. No combination of existing search strategies, then, classifies records with the same accuracy or precision as our final sample.
3. Disentangling Trends in Publication Data
@@
Figure 2.1 indicates that, across measures, the total quantity of potentially-relevant records in PubMed has increased substantially over time, including since 2010.282828Here, we focus on patterns in the production of publications that are similar to those that disclose the results of a clinical trial. We find similar patterns in PubMed as a whole, and in alternative cuts of records. Across fields, the total number of scientific papers, the total number of unique authors listed on scientific papers, the total number of citations given and received by scientific papers, the total number of scientific journals, and the total number of pages in scientific journals exhibit similar patterns. These patterns are consistent with documented increases in the quantity of scientific papers in Bloom et al., (2020) and Park et al., (2023), as well as economy-wide increases in research effort reported in Goldin et al., (2024). In this section, we use our census of clinical trials to decompose this trend. We contrast trends in the production of clinical trials with those documented for other, similar forms of medical research.
There are many potential ways to construct a sample to contrast with data on clinical trial production. Our preferred approach defines this category as the collection of publications that cite clinical trials. This sample includes publications that “look like” clinical trials—a fact that we confirm on inspection—and which necessarily engage with the findings of these studies. Thus, in an imprecise classification task, it is precisely this set of records that are likely to be erroneously included. For example, a keyword search of publication abstracts for “clinical trials” may capture studies that summarize the findings of clinical trials, without reporting new information. Based on our inspection of these records, these types of errors are especially likely for publications in which the writing quality is low, for which it may be challenging—even for a human labeler—to determine whether the record reports novel results or summarizes existing findings.
@@
Alternative methods of constructing a measure of “other” research introduce additional, complicating conceptual problems. Approaches that rely on keyword searches, NLM tags, etc., as Section 2 and its appendices make clear, capture different sets of records over time, as a consequence of changes in database structure. Approaches that examine other publications in the same journals as clinical trials will mechanically suggest shifts in total quantities, if the number of pages in journals change over time or if certain journals become online-only. [^29] Ioannidis et al., (2023) document that the number of medical journals is increasing. In 2024, the Annals of Emergency Medicine, the highest impact journal in the medical sub-field of emergency medicine, shifted to an online-only format. Brainard, (2020) discusses dozens of instances in which medical journals transitioned to online-only formats, then ceased publication entirely.
@@
3.1. Decomposing Trends
Figure 3.1 documents that the quantity of clinical trials—the black line—has remained essentially constant over the past ten years. In contrast, the quantity of other non-trial research—the colored lines—has increased substantially. Observe that we display the percent growth in the quantity of clinical trials in our sample published in each calendar year, alongside the growth in the set of papers that cite clinical trials, constructed in various ways to account for the truncation discussed above. Each series is normalized to begin at zero. This presentation masks substantial differences in the levels of these series. There are 10,903 clinical trials in our census in 2010, and 30,841 publications that cite clinical trials (using three-year citations) in 2013.
@@ 3.2. Clinical Trials as a Productivity Indicator
Stability in public funding suggests, on its own, that the allocation of public funding to clinical trials has not considerably shifted over this time period. One can imagine alternative explanations for this stability. For example, perhaps the share of all projects receiving public funding has remained stable, but the cost of each trial has increased substantially. Alternatively, perhaps public agencies are funding a higher or lower share of studies, but that these changes are perfectly offset by changes in either study success or publication bias.
We view the first potential explanation as unlikely for two reasons. Somewhat suggestively, NIH RePORTER data allow us to tab total expenditures associated with grants acknowledged in these publications. These grant-based expenditure measures, too, are stable over this period. Note, however, that linking research grant dollars to specific scientific papers is an imprecise exercise. See Li, (2017) for an extended discussion of this measurement challenge. More directly, recent work by Sertkaya et al., (2024) suggests that the costs of drug development—measured using the contracts associated with particular study sites in clinical trials—have been essentially constant from 2000 to 2018. The second potential explanation is unlikely from a more conceptual perspective. We find essentially complete stability in this measure over time: any changes in publication bias or trial failure rates must, then, perfectly offset changes in public funding patterns, to yield the flat series here.
Figure 3.3 provides a more granular view of this heterogeneity. We plot quantiles of the distribution of the number of (three-year) citations to publications in our census from publications in leading journals, for each year between 2010 and 2019.323232Appendix D presents these results using alternative cuts of our data and five-year citations. The patterns remain unchanged. Grey areas at the bottom of each plot represent publications that receive zero citations. This distribution is stable over our time period. Observe that we elongate the y-axis to highlight the right-tail of the citation distribution. The quantiles corresponding to the right tail of the citation distribution are constant across years. The left-tail is similarly stable: the share of records receiving zero citations is essentially constant, year-on-year.
@@
We interpret this set of facts as evidence that trends in clinical trial production have been stable over this time period. There are four potential concerns with this interpretation.
First, if the costs of either upstream, basic inputs into clinical trials or the costs of clinical trials themselves have increased substantially over this time, the total resources devoted toward research will have increased, even though Figure 3.1 suggests that counts of trials produced are stable. In our view, this is unlikely to be the case. On the issue of clinical trial costs, the most up-to-date, comprehensive evidence on drug development costs, in our read, is given in Sertkaya et al., (2024), who use internal data from the U.S. Food and Drug Administration and contracts associated with clinical trial sites to compute estimates for the period from 2000 to 2018. In these data, they find that over this period, “the cost of drug development remained relatively stable or may have even decreased.” [^33] This finding revises a frequently cited claim made in popular press interviews (e.g., Herper,, 2022), in academic articles that rely on proprietary data sources (e.g., DiMasi et al.,, 2016), and in surveys on research productivity (Goldin et al.,, 2024)—which suggest that clinical trial and drug development costs are sharply increasing over time. This revision is in the spirit of an argument made by Cockburn, (2006), in response to an earlier debate about drug development costs. Thus, we do not consider the prospect of rising study prices a concern for the interpretation of this pattern. On the issue of basic science inputs into clinical trials, there is little empirical evidence that increases in cost or decreases in the productivity have occurred. A small number of papers examine frictions in markets for basic research. See, for example, Hill and Stein, (2021) and Myers, (2020). To our knowledge, this literature has not provided evidence of substantial shifts in inputs or factor prices over time, which in turn would increase the cost or decrease the productivity of basic science. In fact, changes in the production of basic scientific inputs around the year 2000—enabled by breakthroughs such as the Human Genome Project and the advent of high-throughput screening for drug discovery (see Scannell et al.,, 2012)—suggest, if anything, that costs of basic science inputs have fallen. Moreover, it is reasonable that such trends would affect trends in trial quantity, quality, composition, or cost. Given the absence of such patterns, it is unlikely that changes in basic science threaten the interpretation of our findings. [^34] Constructing a comprehensive measure of medical research productivity, which grapples with these various input markets, remains an open question. To our knowledge, the work that is closest in spirit is Berndt and Cockburn, (2013), who aim to develop a price index for clinical trial research, with the eventual goal of developing price indices for biomedical research more generally.
Second, one might be concerned about how to interpret the x-axis in our plots. In particular, the relationship between clinical trial completion dates and clinical trial start dates varies systematically within and across disease classes (Budish et al.,, 2015). Here, we follow existing literatures that use dates affixed to patents and publications as measures of the timing of research. Practically speaking, as Footnote 14 suggests, records of the start dates of clinical trials are back-filled in administrative databases and, thus, are available only with a considerable lag. We focus on the date of publication, then, for ease of interpretation.
Third, our findings are threatened by publication bias. The existence of publication bias means that counts of scientific publications only partially measure the quantity of research (Andrews and Kasy,, 2019). Our main finding is that the quantity of clinical trials is stable over time. If a trend were in fact present, and our measure were confounded by publication bias, then it would need to be the case that the trend was perfectly offset by the bias. It is reasonable to view cancellation, in this way, as unlikely.
@@
Fourth, and finally, we must assume that the number of publications associated with each clinical trial is constant across our period of interest. We view this assumption as reasonable for three reasons. First, by 2010, nearly all scientific journals in our sample had adopted rules requiring authors of clinical trial publications to adhere to a standardized format when reporting trial results (Schulz et al.,, 2010). Under these rules, there is little scope for authors to, for example, split results across several publications. Second, when constructing our sample of clinical trials, we have excluded publications that re-analyze data generated by an existing study. Third, more suggestively, we find no evidence of a change in this relationship between the size of NIH grants and the quantity of resulting publications reporting the results of a clinical trial during our period of interest. [^35] We inspect this relationship by collecting records from the NIH RePORTER database, for years 2010 to 2019. Specifically, we collect records of the total award associated with funded projects and the quantity of any publications flagged by our procedure as clinical trials. For the median NIH grant, we find that an award of roughly \$100,000 per clinical trial, and a mean of roughly \$200,000 per trial. This is stable across years of data.
Taken together, we view the trends in our data, when set against existing empirical evidence, as indicating that clinical trial publication has been stable over this time period.